The speed c of ocean wave is proportional to the acceleration due to gravity g, wavelengthλ, and densityρof the wave such that c =kgxλyρz,  where k is a dimensionless constant. What is the correct equation for the speed of the ocean?


1
Expert's answer
2021-01-19T13:42:21-0500

The dimension of the acceleration due to gravity is:


"[g] = \dfrac{m}{s^2}"

The dimention of the wavelength is:


"[\lambda] = m"

The dimention of density is:


"[\rho] = \dfrac{kg}{m^3}"

The dimention of the speed is:


"[c] = \dfrac{m}{s}"

Subsituting these values into the formula (and omitting "k" as a dimensionless constant), obtain:


"\dfrac{m}{s} = \dfrac{m^x\cdot m^y\cdot kg^z}{s^{2x}\cdot m^{3z}}"

Comparing the right and left hand side, obtain:


"z = 0\\
2x = 1\\
x+y =1"

Solving this system, obtain:


"x = 0.5\\
y = 0.5"

Then the equation is:


"c = k\sqrt{g\lambda}"

Answer. "c = k\sqrt{g\lambda}" .